What volume of 0.53 mol per DM cube of Potassium Hydroxide would yield 8.2 gram of potassium hydroxide in one vaporization to dryness


1
Expert's answer
2021-05-31T02:24:44-0400

CM = n/V

V = n / CM

n = m / M

M (KOH) = 56.1 g/mol

n (KOH) = 8.2 / 56.1 = 0.15 mol

V (KOH) = 0.15 / 0.53 = 0.28 L = 0.28 dm3